# TSF-Prompt: Integrated Heterogeneous Prompts for Unlocking LLMs in Time Series Forecasting

### v2

@@ Conclusion
This paper proposes TSF-Prompt to unlock the potential of LLMs in time series forecasting. The method incorporates a dual-path prompting mechanism, where the synergistic integration of dynamic selection of learnable soft prompts and semantic mapping of textualized hard prompts enhances LLMs’ comprehension of forecasting tasks. To further improve performance, a cross-modal data fusion architecture is designed through multi-head cross-attention modules, achieving alignment of textual and temporal features. Extensive experiments on multiple real-world datasets validate the effectiveness of TSF-Prompt in time series forecasting.
Despite initial progress, we observe that sole reliance on prompt engineering exhibits significant limitations in effectively activating LLMs’ forecasting capabilities—or, more critically, the intrinsic forecasting capability of LLMs remains unproven. We argue that this fundamental limitation stems from the text-centric training paradigm of LLMs, which inherently struggles to adapt to numerical tasks. Our future work will explore developing foundation models specifically tailored for forecasting tasks, or further demonstrate the predictive capabilities of LLMs through techniques such as chain-of-thought reasoning, fine-tuning, and reinforcement learning.

### v3

@@ Conclusion
This paper proposes Time-Prompt to unlock the potential of LLMs in time series forecasting. This method employs a dual-path prompting mechanism. It utilizes soft prompts to guide the reasoning patterns of the LLM and enhances representations with hard prompts and time series reprogramming. Subsequently, it bridges the modality gap between time series and text through semantic space embedding and cross-modal alignment. Finally, the LLM’s capability for processing continuous values is enhanced through LoRA fine-tuning. Extensive experiments on multiple real-world datasets validate the effectiveness of Time-Prompt in time series forecasting.
Despite preliminary advancements, we find that the time series forecasting capabilities of LLMs remain inadequate. Whether the discrete reasoning abilities of language models can be effective for time series forecasting is still an open question. We argue that this fundamental limitation stems from the text-centric training paradigm of LLMs, which inherently struggles to adapt to forecasting tasks. Moreover, according to the experiments in this paper, we find that small-scale LLMs are more effective than LLMs, and fine-tuning is more effective than prompt engineering in activating predictive capabilities—especially in few-shot scenarios. So our future work will be dedicated to developing time series foundation models or exploring the time-series forecasting capabilities of small-scale LLMs.